Find the critical numbers of
f(x)=|x3−3x2+2|
|Enter the points in increasing order and enter N into any blank you don't need to use.
1
Expert's answer
2019-03-20T09:55:57-0400
"f(x)=|x^3-3x^2+2|""f'(x)=(x^3-3x^2+2)*(x^3-3x^2+2)'\/|x^3-3x+2|"

"f'(x)=(x^3-3x^2+2)*3x*(x-2)\/|x^3-3x^2+2|"



Critical numbers are number where f'(x)=0 or not exist


"(x^3-3x^2+2)*3x*(x-2)\/|x^3-3x^2+2|=0"

"x=0""x=2"

f'(x) not exist at:


"x^3-3x^2+2=0"

"x^3-3x^2+2=(x-1)(x^2-2x-2)=0"

"x=1"

"x^2-2x-2=0"

"D=4-4*(-2)=12"

"x=(2\pm\sqrt{12})/2=1\pm\sqrt{3}""1+\sqrt{3}\approx2.7"

"1-\sqrt{3}\approx-0.7"

So, the numbers in increasing order:


"1-\sqrt{3}; 0; 1; 2; 1+\sqrt{3}"